# A Paradigm Shift to Assembly-like Finite Element Model Updating

### v1

@@ Abstract
[label3]organization=School of Engineering, UWE Bristol,addressline=Frenchay Campus, Coldharbour Lane, city=Bristol, postcode=BS16 1QY, state=England, country=UK
@@
Assembly-like finite element model updating is proposed and validated on a wing model;
The correlation between the finite element model and the experimental results is improved by 20%;
The proposed approach halves the computational cost and offers comparable precision to standard approaches.
@@ 1 Introduction
Over the last three decades, increasing computational power has allowed for the rapid development of finite element model updating (FEMU) methods [1]. This is because a general mismatch between the finite element models (FEMs) and the real systems they are meant to describe usually exists.
@@
The FEMU task is important across all fields of engineering, finding applications such as damage detection [9] and identification [10]. FEMU is also crucial in aeronautics, where having a reliable FEM is pivotal in meeting certification requirements [11]. Further challenges have recently arisen for FEMU in aeronautics, in particular for aircraft wings, because the design paradigm is shifting towards lightweight materials and slenderer wings, in order to improve their aerodynamic efficiency [12]. Hence, developing accurate models is not only necessary for analysing aeroelastic effects, but also for design and controls. In particular, [13] develop a FEM of a scaled aircraft model to be updated with ground vibration testing (GVT) data from different loaded scenarios to estimate flutter onset speed. In [14] a FEMU technique is implemented for positioning a wing box composite material layers for passive aeroelastic suppression, while in [15] model updating is used for enhancing the handling qualities of an aircraft with flexible wings. FEMU has also been implemented for other applications, such as reverse engineering of a fighter aircraft internal structure [16]. Nevertheless, FEMU, in aeronautics, is not only carried out on wings, but also on the full aircraft, e.g. [17], and components, e.g. [18].
@@
Having an accurate structural model of an aircraft structure is not fundamental only for the structural assessment per se, but also for the general life-cycle assessment of the aircraft. Recent advances in the digital twin paradigm have shown that this is vital for building reliable models [26]. These challenges are more challenging nowadays as wings become more flexible. For example, [27, 28] proposes a method for the numerical extraction of modal data from deformed wings (vertical displacement of the wing); however, this was not further explored experimentally. In [29] a component-based FEMU strategy is applied to a composite wing aircraft. Instead of updating the complete structure, the updating is carried out for an individual component, before assembly into the whole system; however, no comparison is given to one-shot approaches, neither in terms of performance nor precision. The latter is addressed, on the same system, in [30], showing favourable results, computationally speaking, for the progressive approach. However, the implementation suggested by Zhao in the aforementioned works, in particular their intermediate steps, is more similar to a hybrid testing approach [31, 32], as usually defined within Civil Engineering, where only part of a structure, or system (e.g. the port wing), is known, rather than to an assembly-like procedure. In fact, EMA data is available only for the wing and for the full assembly, rather than for all the components. Notably, a similar application is carried out for the structural health monitoring of a suspended bridge model in [33]. Nevertheless, from these works it became apparent that the possibility of carrying out assembly-like model assembly, or updating, is appealing as one can consider smaller, in terms of computational loads (e.g. nodes and elements), models, rather than larger and more complex ones. However, a more dedicated approach to the paradigm shift from one-shot and partial model approaches to progressive assembly-like approaches is needed, since these works have not addressed the stepwise characterisation of all subassemblies.
@@
Given the aforementioned motivation, the novelty of this work is to propose and assess a framework for the assembly-like FEMU against the classical direct approach to show its goodness and efficiency. The framework will be tested on a flexible wing model, of which thorough experimental data, involving the full wing and its parts and sub-assemblies, is available in [34, 35]. The hypothesis is that there could be an improvement in model accuracy and computational performance compared to traditional one-shot approaches. The model updating procedure followed in this work is based on a surrogate-based iterative method based on rEGO, which aims at minimising the Modified Total Modal Assurance Criterion (MTMAC) [36] to match the discretised FEM modal response to experimental results. The rEGO is preferred over classical EGO as it provides superior computational efficiency, robustness, and precision [3]. Nevertheless, any other iterative FEMU technique could be implemented with the same concept.
The expected contribution of the work is to prove that the proposed method is more computationally efficient when compared to the classical approach. Thus, the aim of this work is threefold:
@@
    Employ the rEGO-based FEMU technique on a real complex structure;
  3. (iii)
    Obtain an updated discretised FEM of the structure.
@@ 3 The Flexible Wing Model
The reader interested a more profound review of the BeaRDS project is referred to [42, 43, 44, 45, 47] and further information on the XB-2 wing can be found in [43, 45, 34, 53].
@@ 3.2 Experimental Setup
The reader interested in the full experimental work, including details of the setup, acquisition system and post-processing, is referred to [34, 53] for further experimental work on XB-2.
@@ 4 Finite Element Model Updating of the Flexible Wing
First, a baseline discredited FEM is defined as an input file for Ansys Mechanical APDL. The input file takes some parameters, linked to the model properties. The MATLAB implementation finds the set of parameters from rEGO and modifies the input file accordingly, then, still within MATLAB, ANSYS Mechanical APDL is called to run the input file. After this is completed, another file, with the modal parameters results from the FEM, is generated from ANSYS and imported into MATLAB to calculate the MTMAC and recompute the surrogate model. This is shown in Figure 3(a), while, for a practical implementation, the interested reader can refer to the Data Availability Statement.
@@ 4.4 The XB-2 wing
In order to give equal consideration to $\omega_{n}$ and $\bm{\phi}_{n}$ the MTMAC is addressed. The starting MTMAC for the XB-2 wing is 0.30, but the updated models are able to reduce to 0.24 (bottom-up) and 0.23 (top-down). In terms of percentages, the MTMAC of the bottom-up approach improved by 20%, while the top-down approach improved by 23%. Similar results in terms of precision are delivered by the bottom-up and 2 updated FEMs; however, the use of computational power needs to be taken into account. The top-down optimisation over 15 parameters took 1462 evaluations, all on the XB-2 wing model, excluding any top-down iterations for the torque box or the spar models alone. On the other hand to obtain a relevant number the evaluations of all preceding models need to be accounted for. The spar model took 327 evaluations to converge, the torque box model converged after 780 evaluations and, lastly, the final update took 57 evaluations, totalling 1164. It should be noted that the most computationally expensive model (the full wing) is evaluated only 54 times in the bottom-up approach, while, by definition, it is evaluated 1462 in the top-down approach. This shows that the newly proposed bottom-up approach has a similar precision to the traditional approach with a lower computational burden.
@@ 5 Conclusions
    It is the first updated finite element model available in the literature for the benchmark wing

### v2

@@ Abstract
[label3]organization=School of Engineering, University of the West of England,addressline=Frenchay Campus, Coldharbour Lane, city=Bristol, postcode=BS16 1QY, state=England, country=UK
@@
Assembly-like FEM updating method is proposed for structural analysis;
The method is validated using a realistic wing structure model;
FEM-to-experiment correlation improves by 20% using the proposed approach;
Iterative model evaluations are reduced by 50% compared to standard methods;
Accuracy remains comparable to conventional FEM updating.
@@ 1 Introduction
Over the last three decades, increasing computational power has allowed for the rapid development of finite element model updating (FEMU) methods [1]. This is because a general mismatch between the finite element models (FEMs) and the real systems they are meant to describe usually exists [2].
@@
The FEMU task is important across all fields of engineering, finding applications such as damage detection [10], identification [11] and, from a different perspective, in topology optimisation [12]. FEMU is also crucial in aeronautics, where having a reliable FEM is pivotal in meeting certification requirements [13]. Further challenges have recently arisen for FEMU in aeronautics, in particular for aircraft wings, because the design paradigm is shifting towards lightweight materials and slenderer wings, in order to improve their aerodynamic efficiency [14]. Hence, developing accurate models is not only necessary for analysing aeroelastic effects, but also for design and controls. In particular, [15] develop a FEM of a scaled aircraft model to be updated with ground vibration testing (GVT) data from different loaded scenarios to estimate flutter onset speed. In [16] a FEMU technique is implemented for positioning a wing box composite material layers for passive aeroelastic suppression, while in [17] model updating is used for enhancing the handling qualities of an aircraft with flexible wings. FEMU has also been implemented for other applications, such as reverse engineering of a fighter aircraft internal structure [18]. Nevertheless, FEMU, in aeronautics, is not only carried out on wings, but also on the full aircraft, e.g. [19], and components, e.g. [20].
@@
Having an accurate structural model of an aircraft structure is not fundamental only for the structural assessment per se, but also for the general life-cycle assessment of the aircraft. Recent advances in the digital twin paradigm have shown that this is vital for building reliable models [28]. These challenges are more challenging nowadays as wings become more flexible. For example, [29, 30] proposes a method for the numerical extraction of modal data from deformed wings (vertical displacement of the wing); however, this was not further explored experimentally. In [31] a component-based FEMU strategy is applied to a composite wing aircraft. Instead of updating the complete structure, the updating is carried out for an individual component, before assembly into the whole system; however, no comparison is given to one-shot approaches, neither in terms of performance nor precision. The latter is addressed, on the same system, in [32], showing favourable results, computationally speaking, for the progressive approach. However, the implementation suggested by Zhao in the aforementioned works, in particular their intermediate steps, is more similar to a hybrid testing approach [33, 34], as usually defined within Civil Engineering, where only part of a structure, or system (e.g. the port wing), is known, rather than to an assembly-like procedure. In fact, EMA data is available only for the wing and for the full assembly, rather than for all the components. Notably, a similar application is carried out for the structural health monitoring of a suspended bridge model in [35]. Moreover, a substructure-based finite element model updating approach is introduced in [36], wherein selected eigenmodes and derivatives from independent substructures are assembled into a reduced eigenequation, allowing global eigensolutions to be recovered with re-analysis restricted to the relevant substructures. Nevertheless, from these works it became apparent that the possibility of carrying out assembly-like model assembly, or updating, is appealing as one can consider smaller, in terms of computational loads (e.g. nodes and elements), models, rather than larger and more complex ones. However, a more dedicated approach to the paradigm shift from one-shot and partial model approaches to progressive assembly-like approaches is needed, since these works have not addressed the stepwise characterisation of all subassemblies.
@@
Given the aforementioned motivation, the novelty of this work is to propose and assess a framework for the assembly-like FEMU against the classical direct approach to show its goodness and efficiency. The framework will be tested on a flexible wing model, of which thorough experimental data, involving the full wing and its parts and sub-assemblies, is available in [37, 38]. The hypothesis is that there could be an improvement in model accuracy and computational performance compared to traditional one-shot approaches. The model updating procedure followed in this work is based on a surrogate-based iterative method based on rEGO, which aims at minimising the Modified Total Modal Assurance Criterion (MTMAC) [39] to match the discretised FEM modal response to experimental results. The rEGO is preferred over classical EGO as it provides superior computational efficiency, robustness, and precision [4]. Nevertheless, any other iterative optimiser and/or valid goal function could be implemented with the same concept.
The expected contribution of the work is to prove that the proposed, assembly-like FEMU, method is more computationally efficient when compared to the classical approach. Thus, the aims of this work are the following:
@@
    Obtain an updated discretised FEM of a complex structure.
@@ 3 The Flexible Wing Model
The reader interested a more profound review of the BeaRDS project is referred to [45, 46, 47, 48, 49] and further information on the XB-2 wing can be found in [48, 37].
@@ 3.2 Experimental Setup
The reader interested in the full experimental work, including details of the setup, acquisition system and post-processing, is referred to [37, 59]. For further experimental work on the nonlinear dynamics of XB-2, readers are directed to [60].
@@ 4 Finite Element Model Updating of the Flexible Wing
First, a baseline discredited FEM is defined as an input file for Ansys Mechanical APDL. The input file takes some parameters, linked to the model properties. The MATLAB implementation finds the set of parameters from rEGO and modifies the input file accordingly, then, still within MATLAB, ANSYS Mechanical APDL is called to run the input file. After this is completed, another file, with the modal parameters results from the FEM, is generated from ANSYS and imported into MATLAB to calculate the MTMAC and recompute the surrogate model. This is shown in Figure 3(a), while, for a practical implementation, the interested reader can refer to the Data Availability Statement, where the full code implementation of this work is shared. However, it is important to note that for the scope of the validation of the proposed assembly-like FEMU method, any iterative optimiser of choice could replace rEGO in the red dashed-lined box in Figure 3(a).
@@ 4.4 The XB-2 wing
In order to give equal consideration to $\omega_{n}$ and $\bm{\phi}_{n}$ the MTMAC is addressed. The starting MTMAC for the XB-2 wing is 0.30, but the updated models are able to reduce to 0.24 (bottom-up) and 0.23 (top-down). In terms of percentages, the MTMAC of the bottom-up approach improved by 20%, while the top-down approach improved by 23%. Similar results in terms of precision are delivered by the bottom-up and 2 updated FEMs; however, the use of computational power needs to be taken into account. The top-down optimisation over 15 parameters took 1462 evaluations, all on the XB-2 wing model. On the other hand to obtain a relevant number for the bottom-up approach, the evaluations of all preceding models need to be accounted for. The spar model took 327 evaluations to converge, the torque box model converged after 780 evaluations and, lastly, the final update took 57 evaluations, totalling 1164. It should be noted that the most computationally expensive model (the full wing) is evaluated only 54 times in the bottom-up approach, while, by definition, it is evaluated 1462 in the top-down approach. This shows that the newly proposed bottom-up approach has a similar precision to the traditional approach with a lower computational burden, in terms of evaluation and model complexity.
@@ 5 Conclusions
    It is the first updated finite element model available in the literature for the benchmark wing.